OpenStudy (anonymous):

If f(x) = x + 3, find the value of x if f(x) = 9.

OpenStudy (australopithecus):

f(x) is just another way of writing y, it is notation that allows us to discriminate between different functions easier, and allows us to talk about specific points in the function with great ease. in this case we have f(x) = 9 so its just like y = 9 y = x + 3 is the same as f(x) = x + 3 thus 9 = x + 3

OpenStudy (australopithecus):

Hope this explanation helps you better understand the notation

OpenStudy (anonymous):

it does thank u

OpenStudy (australopithecus):

note that you can write (x,y) or (x,f(x)) or (3,f(3))